OpenStudy (anonymous):

exPand this EquatioN for me If you can do So (2x+3)(3x+2)

OpenStudy (anonymous):

\[6x^{2}+13x+5\]

OpenStudy (anonymous):

The above answer I got from \[2x \times 3x = 6x^2, 3x \times 3= 9x, 2x \times 2 = 4x, 2 \times 3 = 6\] and add them all together

OpenStudy (anonymous):

\[1.1255^1/12\] can anybody solve this equation

OpenStudy (anonymous):

1(2x+3)1(3x+2)

OpenStudy (anonymous):

.... penis? You encoded penis in your question with capital letters.